OpenStudy (anonymous):

At intermission, Xavier sold refreshments. Of the items he had, 1/5 were granola bars and 1/2 the remaining items were containers of popcorn. The rest were divided equally among apples, rice cakes, and containers of yogurt. Xavier sold the 8 apples he had. How many items did he originally have?

OpenStudy (texaschic101):

1/5 were granola 1/2 of the remaining. The remaining is 4/5. So 1/2 of 4/5 = (1/2)(4/5) = 4/10 = 2/5 So 1/5 is granola, and 2/5 were containers of popcorn leaving 2/5 to be divided 3 equal ways among apples, rice cakes, and yogurt. So if there was 8 apples, 8 rice cakes, and 8 yogurts that make up 2/5 of this....so 2/5 is 24 so...(2/5) / 24 = (5/5) / x --- 2/5 to 24 = 5/5 to x) cross multiply 2/5x = 24(5/5) 2/5x = 24 x = 24/(2/5) x = 24 * 5/2 x = 120/2 x = 60 check... total 60 1/5 granola : 1/5(60) = 60/5 = 12 2/5 popcorn : 2/5(60) = 120/5 = 24 2/5 (8 apples, 8 rice cakes, 8 yogurts) = 24 ---------------------------------------add 5/5 = 24 + 24 + 12 = 60 He originally had 60 items